The capital asset pricing model is given by the following equation where all terms actually are understood to be expectations:

(Return on asset A = Risk free rate + Beta of A*(Return on the Market - Risk Free rate).

If the Risk Free rate is 5% and the market return is 10%, what is the return on ABC, XYZ, and MNO given the following:

ABC= Beta 1.1

XYZ= beta .80

MNO= beta 1.9
Answer #1

As per Capital Asset Pricing Model (CAPM) Re = Rf + (Rm-Rf) B Where Re = Expected Return or required return of stock Rf = Ris

Calculation of Expected Return of XYZ Rf = 5% Rm - Rf = 10%-5% = 5% B = 0.80 Using the formula Re = Rf + (Rm-Rf) B Re = 5 + 5